A shopkeeper bought two phones for Rs. 8,000 each. After selling the phones, there was a loss of 4% on the 1st phone while a profit of 8% on the 2nd phone. Calculate the overall gain or loss per cent on the whole transaction.

 

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

As the shopkeeper bought both phones at Rs. 8000 each.

Total cost price = Rs. 16,000

Assume that the cost price of the 1st phone is Rs. 100

Consider the deal of phone 1,

As it is given, there is 4% loss, the selling price will be = Rs. 96

For CP = 100, SP = 96

So, for CP = 1, SP = 96/100

Now, given, CP = 8000

Hence, SP = 96/100 × 8000 = 7680

so the selling price of 1st phone = Rs. 7680

Consider the deal of phone 2, there is an 8% profit.

Hence, the selling price will be = Rs. 108

For CP = 100, SP = 108

So, for CP = 1, SP = 108/100

Now, given CP = 8000

hence, SP = 108/100 × 8000 = 8640

Thus, the selling price of 2nd phone = Rs. 8640

Here, the total selling price will be = Rs. 7680 + Rs. 8640 = 16320

Now, it can be seen that,

Total selling price > total cost price i.e. Rs. 16320 > Rs. 16000

, there is a profit of Rs. (16320 – 16000) = Rs. 320

the overall profit percentage will be-

Profit% = (Profit/Total Cost Price) × 100 = (320/16000) × 100 = 2

Thus, there is a total of 2% profit for the whole transaction.

Thanks
Kunal vats
one year ago
Answer
 
Math Primary School
 
 
Expert-Verified Answer
 
Given:
 
A shopkeeper bought two phones for Rs. 8000 each. He sells the first phone at a loss of 4% and also sells the second profit of 8%.
 
To Find:
 
Calculate the overall gain percentage or loss percentage.
 
Solution:
 
1. The cost of each phone is Rs. 8000, Since the shopkeeper has purchased two phones, the cost of two phones is,
 
=> Cost of two mobiles = 2 x 8000 = Rs. 16000.
 
2. Since the first mobile is sold at a loss of 4%,
 
Price at which the first mobile is sold = Cost price - (loss percentage) x Cost price, => Selling price of the first mobile = 8000 - (4/100)x8000,
 
=> Selling price of the first mobile = 8000 - 320,
 
=>Selling price of first mobile = Rs. 7680.
 
3. Since the second mobile is sold at a profit of 8%,
 
=> Price at which the first mobile is sold = Cost price + (profit percentage) x Cost price,
 
=> Selling price of the first mobile = 8000 +
 
(8/100)x8000,
 
=> Selling price of the first mobile = 8000 + 640,
 
=> Selling price of first mobile = Rs. 8640.
 
3. Total selling price of two mobiles = 8640 + 7680 = 16320.
 
>> As the Total selling price is greater than the total cost price, the shopkeeper has gained a profit. ⇒ Profit percentage = (Profit x 100)/Cost price,
 
=>Profit percentage = (320 x 100)/16000,
 
Profit percentage = 2%.